How do you convert 0.13 into a fraction and percent?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Hint: In the above question, we were asked to convert 0.13 into fraction and percent. At first, we will convert 0.13 into a fraction and then we will convert it into a percentage. So, let us see how we can solve this problem. Complete step by step solution: The number of digits after the decimal point is equal to the total number of 0’s. We have 0.13 so it will be 13100 in fraction. As 13100 is already in the lowest form so it will be the answer. For converting 0.13 into percentage we will multiply it with 100 because a percentage is always calculated against 100. So, (0.13 x 100)% = 13% Therefore, 0.13 is 13100 in fraction and 13% in percentage. Additional Information: There are 6 types of fractions; Proper fraction, Improper fraction, Mixed fraction, Like fraction, Unlike fraction, and Equivalent fraction. In a proper fraction, the numerator is smaller and the denominator is larger like 12 . An improper fraction has a larger numerator and smaller denominator like 32 . The fractions which have the same denominator is like fraction for example 12,32,52,72 . 12+32+52+72=(1+3+5+7)2=162=8 . The fractions which have unequal denominators are called unlike fractions for example 12,13 . 12+13=3+26=56 . Equivalent fractions have the same result after simplification, for example, 12and24 are equivalent. Note: Percent simply means a specified amount for every hundred. A decimal number has 2 parts; one is the whole number and the other is the fractional part which is separated by a decimal point. Also, the number of digits after the decimal point is equal to the total number of 0’s.